OpenStudy (anonymous):

what is x in x^2-sinx=0 ?

OpenStudy (anonymous):

can this be answered?

OpenStudy (anonymous):

yeah it has two answers here it is check out http://www.wolframalpha.com/input/?i=x%5E2-sinx%3D0

OpenStudy (anonymous):

one root is x=0 as sin 0 = 0

OpenStudy (anonymous):

jimmrep i think ur wrong cause u cannot neither factor x and nor u can equal them to zero

OpenStudy (anonymous):

so it has no answer?

OpenStudy (anonymous):

Jimmyrep is right. Plug in zero, it is a solution.

OpenStudy (anonymous):

yup = retricesimple as that

OpenStudy (anonymous):

chakaron. It has to have an answer. Lol. http://www.wolframalpha.com/input/?i=x^2-sin(x)%3D0 It has two. As stated in the fundamental theorem of algebra. An n order polynomial is guaranteed to have n solutions.

OpenStudy (anonymous):

it is order 2 - so 2 solutions

OpenStudy (anonymous):

oh. right. thanks.=)

OpenStudy (anonymous):

uhm. is there a step by step solution for this? i don't get why the other x has a value?..

OpenStudy (anonymous):

Consider this: \[x^2-\sin(x)=0 \implies x=\pm \sqrt{\sin(x)}\] Graph y=x, y=sqrt(sin(x)),y=-sqrt(sin(x)) You'll find two intersection points. Those are the solutions. Just pay attention to the top graph. The second one won't have any meaning for you. http://www.wolframalpha.com/input/?i=plot+y%3Dx%2C+y%3Dsqrt%28sin%28x%29%29%2Cy%3D-sqrt%28sin%28x%29%29

OpenStudy (anonymous):

ok. thanks. but still I'm confused with the main problem: y=x^2 and y=sinx. Find the area.